Let’s look at GDP and foreign exchange in Iceland. We can extract the following data from the World Bank, which maintains a database of economic data for all countries:

Fill in the blanks in the table using the formulas from the assigned reading (a bit of algebra is required):

2010

2011

2012

2013

2014

2015

2016

Nominal GDP ($ billions)

1627.11

1708.31

1787.68

1899.68

2020.55

GDP deflator

100.00

102.97

106.35

119.57

122.10

Real GDP ($ billions)

1753.35

1791.93

1869.19

2009.03

Population (total)

318041

319014

320716

323764

327386

330815

335439

Official exchange rate

(Krona per US$)

122.24

115.95

125.08

122.18

116.77

131.92

120.81

Answer these questions:

What was real GDP per capita in 2015? ___________________.

What was real GDP per capita in 2016? ___________________.

Using your answers above, compute the real GDP per capita growth rate for 2016: ___________.

In 2012, your instructor went to Iceland for an academic conference. The Icelandic currency is called the Krona. While there, he spent:

62,540 krona on hotel

10,630 krona on food

15,635 krona on souvenirs

How much did he spend in US dollars? __________________.
Answer #1

(a)

Real GDP in 2015 = $1,869.19 billion

Population in 2015 = 330815

Calculate the real GDP per capita in 2015 -

Real GDP per capita in 2015 = Real GDP in 2015/Population in 2015

Real GDP per capita in 2015 = $1,869.19 billion/330815 = $5,650,257.69

The real GDP per capita in 2015 is $5,650,257.69

(b)

Real GDP in 2016 = $2,009.03 billion

Population in 2016 = 335439

Calculate the real GDP per capita in 2016 -

Real GDP per capita in 2016 = Real GDP in 2016/Population in 2016

Real GDP per capita in 2016 = $2,009.03 billion/335439 = $5,989,255.87

The real GDP per capita in 2016 is $5,989,255.87

(c)

Calculate the real GDP per capita growth rate for 2016 -

Growth rate = [(Real GDP per capita in 2016 - Real GDP per capita in 2015)/Real GDP per capita in 2015] * 100

Growth rate = [($5,989,255.87 - $5,650,257.69)/$5,650,257.69] * 100 = 5.99% or 6%

The real GDP per capita growth rate for 2016 is 5.99 or 6 percent.

(d)

Total spending (in krona) = 62,540 + 10,630 + 15,635 = 88,805 krona

In 2012,

Exchange rate = 125.08 krona per dollar

Total spending (in dollars) = 88,805/125.08 = $709.98

The amount spent in US dollars is $709.98.